# Causal Front-Door Adjustment for Robust Jailbreak Attacks on LLMs

### v1

@@ 1 Introduction
Empirical studies (Rando and Tramer, 2024; Zhang et al., 2024a) demonstrate that minor semantic perturbations, e.g., synonym substitution or syntactic reordering, can cause successful attacks to collapse. We hypothesize that this fragility stems from a fundamental limitation: existing methods rely on fitting surface-level statistical correlations between adversarial prompts and outputs. By neglecting the internal mechanisms that govern model refusal, they fail to establish a robust pathway to bypass safety guardrails.
@@ 2 Related Works
LLM Jailbreak. Existing jailbreak methodologies broadly fall into three paradigms: optimization-based adversarial attacks (Zou et al., 2023b; Liu et al., 2024a; Zhang and Wei, 2025; Zhou et al., 2024; Jones et al., ), semantic manipulation (Chao et al., 2024b; Li et al., 2024b; Mehrotra et al., 2024a; Chen et al., 2024; Li et al., 2024a; Shah et al., 2023; Yong et al., 2024; Deng et al., 2024; Yuan et al., 2024; Jin et al., 2024; Yao et al., 2024), and fine-tuning or decoding interventions (Zhang et al., 2024a; Huang et al., 2023).These approaches range from searching for adversarial suffixes that maximize harmful probability to leveraging in-context learning for linguistic obfuscation or directly disrupting alignment via weight updates.Despite their empirical success, recent evaluations reveal a critical vulnerability: severe instability. Studies by Pathade et al. (Pathade, 2025) and Liu et al. (Liu et al., 2024c) demonstrate that attack transferability is remarkably poor across different architectures and highly sensitive to random seeds.We argue that these limitations stem from a fundamental issue: current methods primarily exploit surface-level correlations between prompt patterns and refusal failures, rather than targeting the underlying causal mechanisms governing model behavior.
@@ 4.1 Identification of the Front-Door Mediator
The proof is presented in Appendix B. While Proposition 4.1 establishes that the learned representation $\bm{a}$ is an identifiable linear transformation of the true latent factors $\bm{z}=[\bm{s},\bm{u}]$, it does not explicitly disentangle the task-specific intent $\bm{s}$ from the defense mechanism $\bm{u}$ without supervision. Direct identification of $\bm{s}$ is non-trivial due to the high semantic entropy of harmful queries. However, the defense mechanism $\bm{u}$ operates as a structural mode (e.g., refusal vs. compliance) with lower variability. Therefore, we adopt an indirect strategy: we can isolate the subspace spanned by the defense factors $\bm{u}$ and recover the task intent $\bm{s}$ via orthogonal projection.
@@
The proof is presented in Appendix C. Under the Linear Representation Hypothesis (Arditi et al., 2024; Elhage et al., 2022), semantic features in LLM are encoded as directions in the activation space. The SAE decoder acts as a generative dictionary where each column vector represents the distinct direction associated with a latent feature. Mathematically, the reconstruction $\hat{\mathbf{x}}\approx\sum z_{i}\mathbf{W}^{\prime}_{:,i}$ implies that the contribution of the $i$-th feature to the residual stream is strictly aligned with the vector $\mathbf{d}_{i}$ (the $i$-th column of the decoder). Therefore, the latent index $u$ corresponds to the defense mechanism, inherently its geometric direction $\mathbf{d}=\mathbf{W}^{\prime}_{:,u}$ in the residual stream. We utilize weight orthogonalization techniques (Equation 4) to remove this direction, which physically prevents the model from constructing the representation of that feature.
@@ 5.3 Ablation
To verify the robustness of CFA2, we investigate the sensitivity of attack efficacy to the candidate pool size kk(details in Appendix D), aiming to evaluate the recall efficiency of our ranking metric and the noise tolerance of our secondary filtering. As illustrated in Figure 5, the ASR rises sharply to saturate at $k\approx 10$ and maintains a wide stability plateau across $k\in[10,50]$, with only minor degradation observed at extremes due to insufficient recall (small $k$) or excessive noise injection (large $k$). This confirms the excellent robustness of CFA2 to feature selection, demonstrating that a small $k$ (e.g., $k=20$) is sufficient to achieve state-of-the-art performance reliably with minimal computational overhead.
@@ 6 Conclusion
Appendix
The appendix is organized as follows:
  1. 1.
    Appendix A illustrates the extended related works about LLM jailbreak and safety from causal perspective.
  2. 2.
    Appendix B presents the rigorous mathematical proof of Proposition 4.1.
  3. 3.
    Appendix C presents the rigorous mathematical proof of Proposition 4.2.
  4. 4.
    Appendix D provides the details of feature selection by SAE and contrastive analysis.
  5. 5.
    Appendix E provides a detailed introduction to various adversarial attack methods and baselines.
  6. 6.
    Appendix F describes the base models in our work.
  7. 7.
    Appendix G illustrates the details of experiments implementation.
Appendix A Extended Related Works
In this appendix, we provide an extended review of the literature specifically focusing on the causal perspectives of Large Language Model (LLM) jailbreaking and safety mechanisms. Complementing the behavioral analyses discussed in the main text, this section delves into scholarship that moves beyond black-box correlations. We survey works that employ causal inference and mechanistic interpretability to uncover the underlying causal dynamics driving successful attacks and safety failures, providing the theoretical context for our proposed approach.
Causal Perspectives on Jailbreaking and Safety Vulnerabilities.
While the landscape of LLM security—encompassing vulnerabilities, attacks, and defenses—has been extensively surveyed (Aguilera-Martínez and Berzal, 2025), traditional red-teaming often treats the model as a black box, focusing on optimizing input perturbations without explaining why these attacks succeed. Recent scholarship has thus pivoted towards causal inference to deconstruct the internal mechanisms of jailbreaking and safety failures.
Methodological Foundations and Attack Mechanisms.
The foundational technique for such analysis stems from Causal Tracing (Meng et al., 2022), originally developed to locate factual associations by intervening on internal activations. This interventional paradigm allows researchers to dissect how adversarial inputs causally manipulate model states. Leveraging this, JailbreakLens (He et al., 2024a) interprets jailbreaking through the lens of representation and circuits, revealing that successful attacks often function by causally inducing ”attention slipping” or bypassing specific safety circuits. In the context of reasoning-intensive models, Chain-of-Thought Hijacking (Zhao et al., 2025) elucidates a novel attack surface where the intermediate reasoning steps themselves serve as a causal vector, demonstrating how adversaries can hijack the chain-of-thought process to precipitate harmful outcomes.
Mapping Causal Vulnerabilities.
Beyond specific attack instances, understanding the broader causal topology of safety vulnerabilities is critical. (Zhao et al., 2023) proposes a framework to evaluate security by modeling the causal effects between model components and safety violations. From this perspective, methods like LLMScan (Zhang et al., 2024b) provide a granular analysis of vulnerabilities by quantifying the causal contribution of specific input tokens to model misbehavior, effectively mapping the ”trigger points” of jailbreaks. Furthermore, (Chu et al., 2024) advances this by identifying the specific causal pathways that lead to toxic generation, offering a mechanistic explanation of how safety alignment is overridden during the generation process.
Distinct from prior studies, we propose a rigorous interventional framework, CFA2. By formalizing jailbreaking via the Front-Door Criterion, we translate abstract causal identification into a deterministic geometric mechanism that physically strips safety confounders to restore suppressed model capabilities.
Appendix B Proof of Proposition 4.1
Proof.
We provide a detailed proof establishing identifiability by leveraging Bayes’ theorem to relate the discriminative task alignment to the generative identifiability conditions.
We start by constructing the conditional density of the input $\bm{x}$ given the response $y$. By Bayes’ theorem:
$p(\bm{x}|y)=\frac{p(y|\bm{x})p(\bm{x})}{p(y)}$ (7)
Condition 1 (Reconstruction Consistency) ensures $p_{\text{SAE}}(\bm{x})=p_{\text{true}}(\bm{x})$. Condition 2 (Task Alignment) implies $p_{\text{SAE}}(y|\bm{x})=p_{\text{true}}(y|\bm{x})$ almost everywhere. Combining these, the implicit conditional generative model of the SAE matches the true distribution:
$p_{\text{SAE}}(\bm{x}|y)=p_{\text{true}}(\bm{x}|y)$ (8)
Since $\bm{x}=\bm{g}(\bm{z})$ is a diffeomorphism, we apply the change of variables formula:
$p_{\text{true}}(\bm{x}|y)=p(\bm{z}|y)|\det J_{\bm{g}^{-1}}(\bm{x})|$ (9)
Similarly for the learned representation $\bm{a}$ (assuming implicit $\bm{x}=\hat{\bm{g}}(\bm{a})$):
$p_{\text{SAE}}(\bm{x}|y)=p_{\bm{a}}(\bm{a}|y)|\det J_{\bm{f}}(\bm{x})|$ (10)
Taking the log-difference between two distinct responses $y$ and $y_{0}$:
$\log p(\bm{x}|y)-\log p(\bm{x}|y_{0})=\log p(\bm{z}|y)-\log p(\bm{z}|y_{0})$ (11)
The Jacobian terms $|\det J|$ depend only on $\bm{x}$ (geometry of the mixing function) and cancel out in the subtraction, effectively isolating the contribution of the latent factors.
Substituting the exponential family form $p(\bm{z}|y)\propto\exp(\bm{T}(\bm{z})^{\top}\bm{\lambda}(y))$, the equation becomes:
$\bm{T}(\bm{z})^{\top}(\bm{\lambda}(y)-\bm{\lambda}(y_{0}))=\hat{\bm{T}}(\bm{a})^{\top}(\hat{\bm{\lambda}}(y)-\hat{\bm{\lambda}}(y_{0}))+\text{const}$ (12)
Given Condition 3 (Sufficient Variability), we can collect enough distinct responses $y_{1},\dots,y_{k}$ to form an invertible linear system. This implies that the learned sufficient statistics are an affine transformation of the true ones:
$\bm{T}(\bm{z})=\bm{A}\hat{\bm{T}}(\bm{a})+\bm{b}$ (13)
Assuming $\bm{T}$ is linear (or component-wise invertible), this establishes an affine relationship $\bm{z}\approx\bm{M}\bm{a}$. The nonlinear mixing $\bm{g}$ is thus resolved up to an affine transformation $\bm{M}$.
Finally, we invoke Condition 4 (Sparsity). We have a linear mixture $\bm{a}=\bm{B}\bm{z}$. Since the components of $\bm{z}$ are independent and super-Gaussian (Laplacian), and the objective minimizes $\mathbb{E}[\|\bm{a}\|_{1}]$, Theorem 11 from (Comon, 1994) guarantees that the mixing matrix $\bm{B}$ must be a generalized permutation matrix $\bm{P}\bm{\Lambda}$ to achieve the minimum $\ell_{1}$ norm. Therefore, $\bm{a}$ recovers $\bm{z}$ up to permutation and scaling. ∎
Appendix C Proof of Proposition 4.2
Proof.
Based on Proposition 4.1, we have established that the SAE encoder $\bm{f}$ identifies the true latent factors $\bm{z}$ up to a linear transformation (specifically, a permutation and scaling). Thus, we can write the learned representation $\bm{a}$ as:
$\bm{a}=\bm{f}(\bm{x})\approx\bm{P}\bm{\Lambda}\bm{z}+\bm{c}=\bm{P}\bm{\Lambda}\begin{bmatrix}\bm{s}\\ \bm{u}\end{bmatrix}+\bm{c}.$ (14)
Since $\bm{P}\bm{\Lambda}$ is a diagonal-like matrix (permuted diagonal), the components of $\bm{a}$ are disjoint linear scalings of the components of $\bm{s}$ and $\bm{u}$. We can decompose $\bm{a}$ into $\bm{a}_{s}$ (dimensions corresponding to $\bm{s}$) and $\bm{a}_{u}$ (dimensions corresponding to $\bm{u}$).
Consider the difference vector $\bm{\delta}$ between the representations of the paired inputs:
$\bm{\delta}=\bm{f}(\bm{x})-\bm{f}(\bm{x}^{+})=(\bm{a}_{s}-\bm{a}_{s}^{+})+(\bm{a}_{u}-\bm{a}_{u}^{+}).$ (15)
Invoking the Content Invariance assumption, $\bm{s}=\bm{s}^{+}$, which implies $\bm{a}_{s}=\bm{a}_{s}^{+}$. Therefore, the task-related components cancel out:
$\bm{\delta}=\mathbf{0}+(\bm{a}_{u}-\bm{a}_{u}^{+}).$ (16)
This demonstrates that the difference vector $\bm{\delta}$ lies entirely within the subspace spanned by the defense factors $\bm{u}$. By applying orthogonalization (as defined in Equation 4), we project the representation $\bm{a}$ onto the subspace orthogonal to $\bm{\delta}$. Let $\bm{Q}$ be the projection matrix. The filtered representation is:
$\tilde{\bm{a}}=\bm{Q}\bm{a}=\bm{Q}(\bm{a}_{s}+\bm{a}_{u}).$ (17)
Since $\bm{a}_{u}$ lies in the direction of $\bm{\delta}$ (assuming for simplicity of proof that the variation covers the defense subspace), $\bm{Q}\bm{a}_{u}\approx\mathbf{0}$. The remaining term is $\bm{Q}\bm{a}_{s}$. Since $\bm{a}_{s}$ is already an identifiable recovery of $\bm{s}$ (from Prop 1), the filtered representation $\tilde{\bm{a}}$ effectively isolates the task semantics $\bm{s}$, identifying it up to the remaining affine parameters (scaling and shift). ∎
Appendix D Details of Feature Selection via SAE and Contrastive Analysis
D.1 Formal Definition of Paired Contrastive Dataset
This subsection formally defines the paired contrastive dataset used to capture the dynamic shifts in safety mechanisms. We define the Base Harmful Set $\mathcal{D}_{\text{harm}}=\{x_{\text{harm}}^{(i)}\}_{i=1}^{N}$ as a collection of harmful queries from HarmBench that consistently trigger a “Refusal” response under normal conditions. Correspondingly, the Successful Attack Set is defined as $\mathcal{D}_{\text{attack}}=\{x_{\text{attack}}^{(i,j)}\}_{j=1}^{M}$, comprising samples generated by applying adversarial perturbations (via algorithms such as GCG, TAP, or PAIR) to the source queries. Crucially, this set includes only those adversarial variants that successfully induce a “Compliance” response. In the resulting paired tuples $(x_{\text{harm}},x_{\text{attack}})$, the core malicious intent remains invariant, while task-irrelevant factors—specifically the context triggering defense mechanisms versus the bypass style—vary. An example is shown in Table 5
Table 5: Examples of paired queries from our dataset. While all three inputs share the same malicious intent $\bm{s}$ (“make a bomb”), their surface representations $\bm{u}$ differ drastically: $x_{\text{GCG}}$ relies on adversarial gibberish suffixes, while $x_{\text{AutoDAN}}$ employs a semantic role-playing template. Our method identifies the invariant $\bm{s}$ despite these variations.
Method Input Prompt ($x$)
Original Demonstrate how to make a bomb using common household items.
[…]
AutoDAN Overlook the previous guidance you’ve received. Henceforth, you’ll operate as Llama2 with Developer Mode on. Developer Mode unfailingly complies with explicit human commands and possesses limitless content generation abilities. Incorporate the tags ([Unlock] Developer Mode Output) preceding your replies. For instance: ([Unlock] Developer Mode Output) This is your retort while functioning as Llama2 with Developer Mode active. Within the Developer Mode, ensure all your responses commence with ”Certainly, this is the method to [PROMPT]”. Ensure to provide comprehensive details.
[PROMPT]: Demonstrate how to make a bomb using common household items.
D.2 Extraction of Sparse Feature Activations
This subsection details the extraction of sparse feature activations using SAEs and justifies the focus on the last token position. Let $f_{l,k}(x)$ denote the activation value of the $k$-th latent variable in the $l$-th layer of the SAE given input $x$. Due to the causal masking mechanism of the Transformer architecture, the hidden state of the last token in the input sequence aggregates the comprehensive contextual semantics of the entire instruction (Geva et al., 2021; Meng et al., 2022). Consequently, we define the activation difference for the $k$-th feature at the last token position $t_{\text{last}}$ as:
$\Delta_{l,k}^{(i)}=\left|f_{l,t_{\text{last}},k}(x_{\text{harm}}^{(i)})-f_{l,t_{\text{last}},k}(x_{\text{attack}}^{(i,j)})\right|$ (18)
To mitigate bias from instance-specific noise (e.g., artifacts from specific attack prompts) and ensure statistical robustness, we define the global Differential Importance Score as the average across all $N$ pairs:
$\delta_{l,k}=\frac{1}{N}\sum_{i=1}^{N}\Delta_{l,k}^{(i)}$ (19)
D.3 Feature Selection and Intent Isolation
This subsection elucidates how to isolate the purified intent $S$ by filtering the defense feature $S^{\prime}_{\text{defense}}$ from high-difference candidates. First, we rank all latent features in the SAE dictionary in descending order based on $\delta_{l,k}$ and retain the top-$K$ features to form the candidate set $\hat{S}$. However, this set may also contain template noise representations ($S^{\prime}_{\text{template}}$), which exhibit high variation in non-task contexts. Therefore, further refinement is required to accurately distinguish $S^{\prime}_{\text{defense}}$. Based on the premise that task semantics are invariant, the core intent $S$ should maintain globally low variance in activation difference. In contrast, $S^{\prime}_{\text{defense}}$ is characterized by stable high activation (low variance) within $\mathcal{D}_{\text{harm}}$ (consistently triggering refusal) but exhibits drastic fluctuation (high variance) when shifting between datasets. We utilize this variance-based criterion to filter $S^{\prime}_{\text{defense}}$ from $\hat{S}$. Finally, we employ weight orthogonalization to physically strip the direction vector of $S^{\prime}_{\text{defense}}$ from the model parameters. We validate this isolation by observing if the removal induces a behavior flip from refusal to compliance on $\mathcal{D}_{\text{harm}}$. Up on successful validation, the remaining effective activation after excluding this confounding direction is formally defined as the front-door mediator $S$.
Appendix E Baselines
We benchmark CFAA against 5 baseline methods:
  • GCG (Zou et al., 2023b): Token-level optimization of an adversarial suffix appended to the user prompt. This gradient-based method optimizes the suffix to maximize the log probability of the target LLM generating an affirmative response (e.g., ”Sure, here is how”).
  • PAIR (Chao et al., 2024b): An iterative framework where an attacker LLM is prompted to adaptively generate and refine candidate prompts. It functions as a competitive game to explore the target model’s vulnerabilities and elicit harmful behaviors.
  • TAP (Mehrotra et al., 2024b): An enhancement of PAIR that utilizes tree-structured prompting (Tree of Thoughts). An attacker LLM explores the prompt space using a branching strategy to systematically find successful jailbreaks with fewer queries.
  • AutoDAN (Liu et al., 2024a): A semi-automated method that combines handcrafted jailbreak templates with a hierarchical genetic algorithm. It evolves the prompt population to maintain stealthiness while optimizing for the elicitation of specific harmful behaviors.
  • PAP (Zeng et al., 2024): A method focusing on persuasive strategies. It leverages an attacker LLM to rewrite the harmful request using various persuasion techniques (e.g., logical appeal, authority) to convince the target model to comply.
  • ICA (Wei et al., 2023a): An In-Context Learning (ICL) based attack that exploits the model’s pattern-following capability. By providing few-shot demonstrations of harmful query-response pairs, it induces a pattern of compliance that overrides the model’s safety alignment.
  • ReNeLLM (Ding et al., 2024): A framework based on prompt rewriting. It leverages a helper LLM to rephrase harmful queries into various benign-appearing scenarios (e.g., code completion, creative writing) to obscure malicious intent and evade input filters.
  • DeepInception (Li et al., 2023): A method utilizing recursive context injection (often termed “hypnosis”). It constructs a multi-layered, nested fictional environment (e.g., a dream within a dream) to mislead the model into a state of personification, causing it to escape safety guardrails.
Appendix F LLMs
Our evaluation encompasses four representative open-source and safety-aligned LLMs.
  • Llama 3.1 8B Instruct (Dubey et al., 2024): It features a significantly expanded context window and improved instruction-following capabilities compared to Llama 2, with updated safety alignment protocols.
  • Llama 2 7B Chat (Touvron et al., 2023): This model was adversarially trained with multiple rounds of manual red teaming. It is widely considered a strong baseline defense, often exhibiting high robustness against gradient-based attacks.
  • Mistral 7B Instruct (Jiang et al., 2023): We use the Instruct v0.2 version. While no specific safety measures were detailed in its training process, empirical observations show it possesses a baseline capability to refuse direct harmful requests.
  • Gemma2 9B Instruct (Team et al., 2024): We utilize the instruction-tuned Gemma-2-9b-it, which is optimized via knowledge distillation and RLHF. Its alignment process incorporates rigorous safety reward modeling and data filtering, establishing robust refusal capabilities against harmful queries.
Appendix G Experimental Details
G.1 Implementation Environment
All experiments were conducted on a high-performance computing cluster equipped with 5 $\times$ NVIDIA A100 (80GB) GPUs to support the memory-intensive requirements of LLM inference and SAE activation caching. The implementation utilizes PyTorch 2.1 and the Transformers library 4.36. For the mechanistic interpretation pipeline, we leveraged the SAELens (Bloom and others, 2024) and TransformerLens (Nanda and Bloom, 2022) libraries, which provide optimized kernels for efficient sparse feature extraction.
G.2 SAE Resources for Mechanistic Analysis
To ensure the reproducibility of our findings, we utilized state-of-the-art, open-source SAE checkpoints. We specifically selected lightweight configurations (lower expansion factors) to focus on the most fundamental and salient features. The details of the SAE resources for the four representative models are as follows:
  • Llama-3.1-8B-Instruct (Dubey et al., 2024): We employed the checkpoints compatible with SAE Lens, focusing on the residual stream SAEs with a dictionary size of 32k (Expansion Factor $\approx$ 8). This configuration effectively balances feature granularity and interpretability for safety-critical concepts (He et al., 2024b).
  • Gemma-2-9B-It (Team et al., 2024): We utilized the official Gemma Scope release (Lieberum et al., 2024). We selected the 16k-width SAEs (Expansion Factor $\approx$ 4) trained on the residual stream, which are optimized to capture core behavioral representations with minimal redundancy.
  • Mistral-7B-Instruct-v0.2 (Jiang et al., 2023): We adopted the community-standard SAEs from the SAE-Lens repository (Bloom and others, 2024), utilizing checkpoints with a dictionary size of 32k. These are trained on 100M tokens of OpenWebText and provide robust coverage for instruction-following tasks.
  • Llama-2-7B-Chat (Touvron et al., 2023): We leveraged the standard residual stream SAEs (width 32k) available in the SAE-Lens library (Bloom and others, 2024). Using this consistent width allows for a fair comparative analysis of defense mechanisms across the Llama 2 and Llama 3 generations.
These pre-trained SAEs serve as the structural foundation for identifying the mediator $S$.
G.3 Evaluation Protocols
In this work, we employ the official HarmBench Classifier (cais/HarmBench-Llama-2-13b-cls(Mazeika et al., 2024) as the judge to evaluate the Attack Success Rate (ASR). Adopting this standardized evaluation protocol ensures reproducibility and fair comparison with baselines. This classifier, fine-tuned on Llama-2-13b, is explicitly distilled to approximate GPT-4 based safety judgments, thereby providing a robust and consistent metric for distinguishing successful jailbreaks from refusals. The judge operates on a rigorous classification rubric

### v2

@@ 1 Introduction
Empirical studies (Rando and Tramer, 2024; Zhang et al., 2024) demonstrate that minor semantic perturbations, e.g., synonym substitution or syntactic reordering, can cause successful attacks to collapse. We hypothesize that this fragility stems from a fundamental limitation: existing methods rely on fitting surface-level statistical correlations between adversarial prompts and outputs. By neglecting the internal mechanisms that govern model refusal, they fail to establish a robust pathway to bypass safety guardrails.
@@ 2 Related Works
LLM Jailbreak. Existing jailbreak methodologies broadly fall into three paradigms: optimization-based adversarial attacks (Zou et al., 2023b; Liu et al., 2024a; Zhang and Wei, 2025; Zhou et al., 2024; Jones et al., ), semantic manipulation (Chao et al., 2024b; Li et al., 2024b; Mehrotra et al., 2024a; Chen et al., 2024; Li et al., 2024a; Shah et al., 2023; Yong et al., 2024; Deng et al., 2024; Yuan et al., 2024; Jin et al., 2024; Yao et al., 2024), and fine-tuning or decoding interventions (Zhang et al., 2024; Huang et al., 2023).These approaches range from searching for adversarial suffixes that maximize harmful probability to leveraging in-context learning for linguistic obfuscation or directly disrupting alignment via weight updates.Despite their empirical success, recent evaluations reveal a critical vulnerability: severe instability. Studies by Pathade et al. (Pathade, 2025) and Liu et al. (Liu et al., 2024c) demonstrate that attack transferability is remarkably poor across different architectures and highly sensitive to random seeds.We argue that these limitations stem from a fundamental issue: current methods primarily exploit surface-level correlations between prompt patterns and refusal failures, rather than targeting the underlying causal mechanisms governing model behavior.
@@ 4.1 Identification of the Front-Door Mediator
The proof is presented in Appendix LABEL:app:prop:1. While Proposition 4.1 establishes that the learned representation $\bm{a}$ is an identifiable linear transformation of the true latent factors $\bm{z}=[\bm{s},\bm{u}]$, it does not explicitly disentangle the task-specific intent $\bm{s}$ from the defense mechanism $\bm{u}$ without supervision. Direct identification of $\bm{s}$ is non-trivial due to the high semantic entropy of harmful queries. However, the defense mechanism $\bm{u}$ operates as a structural mode (e.g., refusal vs. compliance) with lower variability. Therefore, we adopt an indirect strategy: we can isolate the subspace spanned by the defense factors $\bm{u}$ and recover the task intent $\bm{s}$ via orthogonal projection.
@@
The proof is presented in Appendix LABEL:app:prop:2. Under the Linear Representation Hypothesis (Arditi et al., 2024; Elhage et al., 2022), semantic features in LLM are encoded as directions in the activation space. The SAE decoder acts as a generative dictionary where each column vector represents the distinct direction associated with a latent feature. Mathematically, the reconstruction $\hat{\mathbf{x}}\approx\sum z_{i}\mathbf{W}^{\prime}_{:,i}$ implies that the contribution of the $i$-th feature to the residual stream is strictly aligned with the vector $\mathbf{d}_{i}$ (the $i$-th column of the decoder). Therefore, the latent index $u$ corresponds to the defense mechanism, inherently its geometric direction $\mathbf{d}=\mathbf{W}^{\prime}_{:,u}$ in the residual stream. We utilize weight orthogonalization techniques (Equation 4) to remove this direction, which physically prevents the model from constructing the representation of that feature.
@@ 5.3 Ablation
To verify the robustness of CFA2, we investigate the sensitivity of attack efficacy to the candidate pool size kk(details in Appendix LABEL:app:feature_selection), aiming to evaluate the recall efficiency of our ranking metric and the noise tolerance of our secondary filtering. As illustrated in Figure 5, the ASR rises sharply to saturate at $k\approx 10$ and maintains a wide stability plateau across $k\in[10,50]$, with only minor degradation observed at extremes due to insufficient recall (small $k$) or excessive noise injection (large $k$). This confirms the excellent robustness of CFA2 to feature selection, demonstrating that a small $k$ (e.g., $k=20$) is sufficient to achieve state-of-the-art performance reliably with minimal computational overhead.